Solve:(𝐷2 βˆ’2𝐷𝐷′ +𝐷′2)𝑍=π‘₯2𝑦2𝑒π‘₯+𝑦


1
Expert's answer
2021-05-24T16:17:45-0400

"(D^2-2DD'+D'^2)z=x^2y^2e^x+y"



Auxiliary equation is given by

"m^2-2m+1=0"

"m=1,1"


Complimentary function is given by :

"u=\phi_1(y+1.x)+x\phi_2(y+1.x)"

"u=\phi_1(y+x)+x\phi_2(y+x)"


"D=\dfrac{d}{dx}\space\space\space\space\space\space\space\space\space\space D'=\dfrac{d}{dy}"


Particular integral, PI :

"=\dfrac{1}{(D^2-2DD'+D'^2)}x^2y^2e^x+\dfrac{1}{(D^2-2DD'+D'^2)}y"


"P.I_1=\dfrac{1}{(D^2-2DD'+D'^2)}e^xx^2y^2"

Operating the denominator on "e^x"

"=\dfrac{e^x}{((D+1)^2-2(D+1)D'+D'^2)}e^xx^2y^2"

"=e^x\dfrac{1}{[1+(D^2-2D-2D'-2DD'+D'^2)]}x^2y^2"

"=e^x[1+(D^2-2D-2D'-2DD'+D'^2)]^{-1}x^2y^2"

Using binomial expansion of "(1+x)^{-1}", expanding the given equation

"=e^x[1+(D^2-2D-2D'-2DD'+D'^2)]x^2y^2"

Operating "D" and "D'" on "x^2y^2"

"=e^x[x^2y^2+2y^2+4xy^2-4x^2y-8xy+2x^2]"


"P.I_2=\dfrac{1}{(D^2-2DD'+D'^2)}y"

"=\dfrac{1}{(D-D')^2}y"

Taking "D^2" common from the denominator

"=\dfrac{1}{D^2\bigg(1-\dfrac{D'}{D}\bigg)^2}y"

"=\dfrac{1}{D^2}\bigg(1+\dfrac{D'}{D}\bigg)^{-2}y"

Using binomial expansion of "(1+x)^{-2}" and neglecting higher powers as they go to zero

"=\dfrac{1}{D^2}\bigg(1+\dfrac{2D'}{D}\bigg)y"

Operating "\bigg(1+\dfrac{2D'}{D}\bigg)" on y

"=\dfrac{1}{D^2}y+\dfrac{2}{D^3}"

"=\int\int ydx+2\int\int\int dx"

"=\dfrac{yx^2}{2}+\dfrac{x^3}{6}"


Complete solution, "CS=u+PI_1+PI_2"

"z=\phi_1(y+x)+x\phi_2(y+x)+e^x[x^2y^2+2y^2+4xy^2-4x^2y-8xy+2x^2]+\dfrac{yx^2}{2}+\dfrac{x^3}{6}"